# A mixed-precision quantum-classical algorithm for solving linear systems

### v1

@@ Abstract
Index Terms:
Quantum computing, Linear systems, Quantum Singular Value Transformation, Mixed-precision algorithms, Iterative refinement.
@@ I Introduction
Iterative refinement has been recently studied for the HHL algorithm [36, 39] or for optimization problems [31]. In this paper we consider the LS problem solved via the QSVT method which offers several advantages: the matrix does not need to be Hermitian and even not square (in this case we solve a least squares problem), the QSVT method can exploit efficient block-encoding techniques and appears to be well suited for Large Scale Quantum (LSQ) architectures. However the quantum circuit for the QSVT requires a polynomial approximation that is very expensive if we want to achieve acceptable accuracy (typically better than $10^{-5}$ [32]). This has motivated our choice for developing a mixed-precision algorithm which combines the speed and limited precision of the QSVT with iterative refinement in higher precision. Mixed-precision techniques are widely used in HPC algorithms [1]. They have been initially motivated by the emergence of GPU accelerators that provide high performance when computing in lower precision arithmetic, for instance using tensor core units [4]. For hybrid CPU/GPU architectures we can achieve high performance of the solver while preserving the accuracy of the higher precision and we investigate how similar techniques can be applied with the QSVT method. Similarly to the CPU/GPU case, our iterative refinement is also hybrid since it uses a classical processor (to compute the residual and solution update) and a quantum processor (for the QSVT solver). We adapt the iterative refinement to the case where the LS problem is solved via the QSVT and we present an error and complexity analysis to evaluate the quantum ans classical costs of our algorithm. We also show that our approach provides a significant advantage in complexity compared to directly solving the LS problem with QSVT in higher precision.
@@
[…]
@@ II-A1 Block-encoding of matrices
When we apply $U$ to |$\ket{0}_{a}\ket{\psi}_{d}$ , where |$\ket{0}_{a}$ corresponds to the ancilla qubits and |$\ket{\psi}_{d}$ corresponds to the data qubits, we get
@@ II-A2 QSVT definition
($\Sigma=diag(\sigma_{1},\dots,\sigma_{N})$ and $W,~{}V$ unitary matrices), given a polynomial $P$ of degree $d$, then the QSVT of $A$ using $P$ is expressed by
@@ II-A3 QSVT implementation
This operator can be used to apply the QSVT in a quantum computer, according to the polynomial $P$, to the matrix $A$
@@ II-A4 Solving linear systems with QSVT
In practice we want to find an $\frac{\epsilon}{2\kappa}$-approximation to $\frac{1}{2\kappa}\frac{1}{x}$, where $\kappa$ is the condition number of the matrix to be inverted, and $\epsilon$ the error on $[-1,1]\setminus[\frac{-1}{\kappa},\frac{1}{\kappa}]$. The inverse function is difficult to approximate with a polynomial (not continuous, infinite values in 0…). We need to find an odd function approximating the inverse function on $[-1,1]\setminus[\frac{-1}{\kappa},\frac{1}{\kappa}]$ that would be easier to approximate using a polynomial. A function that can be used is
@@ II-B Mixed-precision iterative refinement for linear systems
This process can then be repeated until we obtain a satisfying $\tilde{x}$. Using the same precision arithmetic throughout the process (fixed-precision iterative refinement) is classically used to improve the accuracy of a computed solution or added to ameliorate a potentially instable solver [19, p. 232]. However, with the emergence of processors that propose much faster computation using lower precision arithmetic, it became attractive to combine different precisions in order to exploit the high performance provided by some processors or computational units (e.g., tensor cores for NVIDIA GPUs), resulting in so-called mixed-precision iterative refinement algorithms.
@@ III-A Algorithm
Suppose we can solve $Ax=b$ via the QSVT method with a (low) accuracy $\epsilon_{l}$, i.e., that can produce a solution $\tilde{x}$ such that $\|{x-\tilde{x}}\|\leq\epsilon_{l}\lVert x\rVert$. Following [28, p. 126], to obtain an accuracy (relative error) of order $\epsilon_{l}$ on the non-normalized solution $\tilde{x}$, we need to approximate the inverse function on $[-1,-1/\kappa]\cup[1/\kappa,1]$ with an error $\epsilon^{\prime}=\mathcal{O}(\epsilon_{l}/\kappa)$.
@@
Input: $A$, $b$, QSVT accuracy $\epsilon_{l}$, targeted accuracy $\epsilon$ at precision $u$.
@@ III-C1 Quantum cost
# solves $1$ $\leq\left\lceil\dfrac{log(\epsilon)}{log(\kappa\;\epsilon_{l})}\right\rceil$
@@ III-C4 Practical example
Solution $\mathcal{O}(4^{n}+\log(1/\epsilon))$
@@
Solution $\mathcal{O}(4^{n}+\log(1/\epsilon))$
@@ IV-B Results
We present in Figure 3 the evolution of the scaled residual obtained at each iteration, until convergence. We consider a small condition number $\kappa=10$ and three values for $\epsilon_{l}$. The targeted accuracy is $\epsilon=10^{-11}$. We observe that in all cases the bound $\lceil\log(\epsilon)/\log(\epsilon_{l}\kappa)\rceil$ obtained in Theorem III.1 is a sharp estimate for the iteration count. We have also observed that due to the small values of $\kappa$ the scaled residual is here also a good estimate of the forward error, as expected from Equation (5).

### v2

@@ I Introduction
Iterative refinement has been recently studied for the HHL algorithm [36, 39] or for optimization problems [31]. In this paper we consider the LS problem solved via the QSVT method which offers several advantages: the matrix does not need to be Hermitian and even not square (in this case we solve a least squares problem), the QSVT method can exploit efficient block-encoding techniques and appears to be well suited for Large Scale Quantum (LSQ) architectures. Note that our work is carried out in an LSQ context and not NISQ (Noisy Intermediate-Scale Quantum) due to the excessive depth of quantum circuits for the QSVT algorithm. The quantum circuit for the QSVT requires a polynomial approximation that is very expensive if we want to achieve acceptable accuracy (typically better than $10^{-5}$ [32]). This has motivated our choice for developing a mixed-precision algorithm which combines the speed and limited precision of the QSVT with iterative refinement in higher precision. Mixed-precision techniques are widely used in HPC algorithms [1]. They have been initially motivated by the emergence of GPU accelerators that provide high performance when computing in lower precision arithmetic, for instance using tensor core units [3]. For hybrid CPU/GPU architectures we can achieve high performance of the solver while preserving the accuracy of the higher precision and we investigate how similar techniques can be applied with the QSVT method. Similarly to the CPU/GPU case, our iterative refinement is also hybrid since it uses a classical processor (to compute the residual and solution update) and a quantum processor (for the QSVT solver). We adapt the iterative refinement to the case where the LS problem is solved via the QSVT and we present an error and complexity analysis to evaluate the quantum ans classical costs of our algorithm. We also show that our approach provides a significant advantage in complexity compared to directly solving the LS problem with QSVT in higher precision.
@@
Notation: Unless otherwise stated, $\|{.}\|$ denotes the Euclidean norm for vectors and the spectral norm for matrices.
@@ II-A1 Block-encoding of matrices
[…]
@@ II-A2 QSVT definition
($\Sigma=diag(\sigma_{1},\dots,\sigma_{N})$ and W,VW,~V unitary matrices), given a polynomial $P$ of degree $d$, then the QSVT of $A$ using $P$ is expressed by
@@ II-A3 QSVT implementation
This operator can be used to apply the QSVT in a quantum computer, according to the polynomial $P$, to the matrix AA
@@ II-A4 Solving linear systems with QSVT
In practice we want to find an $\frac{\epsilon}{2\kappa}$-approximation to $\frac{1}{2\kappa}\frac{1}{x}$, where $\kappa$ is the condition number of the matrix to be inverted, and $\epsilon$ the error on $1,1]\setminus[\frac{-1}{\kappa},\frac{1}{\kappa}]$. The inverse function is difficult to approximate with a polynomial (not continuous, infinite values in 0…). We need to find an odd function approximating the inverse function on $1,1]\setminus[\frac{-1}{\kappa},\frac{1}{\kappa}]$ that would be easier to approximate using a polynomial. A function that can be used is
@@ II-B Mixed-precision iterative refinement for linear systems
This process can then be repeated until we obtain a satisfying $\tilde{x}$. Using the same precision arithmetic throughout the process (fixed-precision iterative refinement) is classically used to improve the accuracy of a computed solution or added to ameliorate a potentially instable solver [20, p. 232]. However, with the emergence of processors that propose much faster computation using lower precision arithmetic, it became attractive to combine different precisions in order to exploit the high performance provided by some processors or computational units (e.g., tensor cores for NVIDIA GPUs), resulting in so-called mixed-precision iterative refinement algorithms.
@@ III-A Algorithm
Suppose we can solve $Ax=b$ via the QSVT method with a (low) accuracy $\epsilon_{l}$, i.e., that can produce a solution $\tilde{x}$ such that $\|{x-\tilde{x}}\|\leq\epsilon_{l}\lVert x\rVert$. Following [28, p. 126], to obtain an accuracy (relative error) of order $\epsilon_{l}$ on the non-normalized solution $\tilde{x}$, we need to approximate the inverse function on $1,-1/\kappa]\cup[1/\kappa,1]$ with an error $\epsilon^{\prime}=\mathcal{O}(\epsilon_{l}/\kappa)$.
@@
Input: AA, $b$, QSVT accuracy $\epsilon_{l}$, targeted accuracy $\epsilon$ at precision $u$.
@@ III-C1 Quantum cost
# solves 11 $\leq\left\lceil\dfrac{log(\epsilon)}{log(\kappa\;\epsilon_{l})}\right\rceil$
@@ III-C4 Practical example
Solution $\mathcal{O}(4^{n}+\log(1/\epsilon))$ -
@@
Solution $\mathcal{O}(4^{n}+\log(1/\epsilon))$ -
@@ IV-B Results
We present in Figure 3 the evolution of the scaled residual obtained at each iteration, until convergence. We consider a small condition number $\kappa=10$ and three values for $\epsilon_{l}$. The targeted accuracy is $\epsilon=10^{-11}$. We observe that in all cases the bound $\lceil\log(\epsilon)/\log(\epsilon_{l}\kappa)\rceil$ obtained in Theorem III.1 is a sharp estimate for the iteration count. We also observe that due to the small values of $\kappa$ the scaled residual is here also a good estimate of the forward error, as expected from Equation (5).
@@ V Conclusion
Acknowledgement
This work is part of the HQI initiative (www.hqi.fr) and is supported by France 2030 under the French National Research Agency award number “ANR-22-PNCQ-0002”.